OpenStudy (anonymous):

(solve for the system of equations) -7x-3y=-1 9x+5y=7

OpenStudy (anonymous):

Matrix? Solve for augment matrix!!

OpenStudy (anonymous):

-7 -3 -1 9 5 7

OpenStudy (anonymous):

im not working with matrix

OpenStudy (anonymous):

Oop sorry then solve by substitution?

OpenStudy (anonymous):

Just solve for x on each one first. Like this: -7x-3y=-1 -7x =-1+3y ---I added 3y to both sides x =1/7-3y/7 ---Divided -7 to both sides.

OpenStudy (anonymous):

And then after you do it for the second equation we can susbtitue x from both equations and then we only have to solve for y

OpenStudy (anonymous):

9x+5y=7 x =7/9-5y/9

OpenStudy (anonymous):

x =1/7-3y/7 x =7/9-5y/9 7/9-5y/9 =1/7-3y/7 Solve for y now. After you find y plug it back in to one of the original equations again and solve for x.

OpenStudy (anonymous):

ok thanks man(: